Pls Include solution:)

 A coin tossed thrice. Find: 

 a. The probability of getting exactly 3 heads. 

 b. The probability of getting exactly 2 tails.

 c. The probability of getting atmost 1 head. 

 d. The probability of getting atleast 1 tail.


1
Expert's answer
2022-05-12T09:03:36-0400

The sample space of a fair coin flip is "\{H, T\}." The sample space of a sequence of three fair coin flips is all "2^3=8" possible sequences of outcomes:


"\{ HHH, HHT, HT H, HT T,""T HH, T HT, T T H, T T T\}"


a.

Exactly 3 heads: "HHH"


"P(exactly\ 3\ heads)=\dfrac{1}{8}"

b.

Exactly 2 tails: "HTT, THT, TTH"


"P(exactly\ 2\ tails )=\dfrac{3}{8}"


c.

At most 1head: "HTT, THT, TTH, TTT"


"P(at\ most\ 1\ head)=\dfrac{4}{8}=\dfrac{1}{2}"

d.

At least 1 tail:"HHT, HT H, HT T, T HH, T HT, T T H, T T T"


"P(at\ least\ 1\ tail)=\dfrac{7}{8}"